You are here

Beam Delivery News

Primary tabs

TItle Date Excerpt Attached images
Beam delivery update 17 October 2014

The cyclotron is presently delivering ~110 µA to BL1A and 50 µA to BL2A; BL2C1 is set up for proton therapy tomorrow morning. High-current delivery to BL1A and BL2A, interrupted each morning for proton therapy, is expected to continue through the weekend.

Beam delivery update: Proton therapy starting tomorrow 16 October 2014

The cyclotron is delivering ~110 µA to BL1A and 50 µA to BL2A. This is expected to continue until the start of proton therapy at 08:00 tomorrow morning. BL2C1 was set up for PT this afternoon.

Beam delivery update 15 October 2014

The cyclotron is now running pretty much normally, delivering ~110 µA to BL1A and 50 µA to BL2A. The STF group hasn’t yet decided if they’re going to request beam overnight or if BL2C will stay off until tomorrow's switchover to 2C1.

Further to last night's post… 14 October 2014

The main magnet is on, the RF has been tested, and inflector/deflector conditioning is underway. The present hope is to inject beam some time later today for full production tomorrow as originally scheduled.

Mini-shutdown update: Cyclotron startup delayed 13 October 2014

A power bump due to generator testing tripped off the vacuum pump on the space between the cyclotron tank seals on Friday, causing a partial vent of the tank. The vacuum is still recovering. It looks like Wednesday is the earliest at which beam might be injected as a result.

Mini-shutdown update 09 October 2014

We’re about half-way through the scheduled mini-shutdown. Injection and startup are scheduled for Tuesday with beam delivery to resume Wednesday.

Beam delivery update 01 October 2014

The cyclotron is delivering ~105 µA to BL1A, 50 µA to BL2A, and ~95 µA to BL2C4. This is expected to continue at least overnight and possibly through the weekend.

Beam delivery update: ISAC maintenance tomorrow 29 September 2014

The cyclotron is currently delivering ~105 µA to BL1A, 5 µA to BL2A (for end-of-target target scans), and ~95 µA to BL2C4. BL2A is expected to go off at some point this evening.

Beam delivery update 20 September 2014

BL1A is the only beamline running right now; it’s taking ~105 µA and will likely do so through the weekend. BL2A is off due to target issues while BL2C is off until tomorrow morning’s proton therapy.

Beam delivery update 18 September 2014

Beam is currently off due to RF problems. The plan is to deliver ~105 µA to BL1A, 50 µA to BL2A, and 95 µA to BL2C4 overnight before inserting the pepper pot for proton therapy at 08:00 tomorrow morning.

Pages

Subscribe to Group News